# Visual Concept-driven Image Generation with Text-to-Image Diffusion Model

### v2

@@ Abstract
Keywords:
Diffusion model Segmentation Concept-driven generation
1 Introduction
@@ 2 Related Work
3 Methodology
Figure 2: Overview of our EM-style Optimization Framework. Unique tokens are first assigned to the concepts. In this example, three tokens are assigned: [v1] to the male subject in the first image, [v2] to the female subject in the second image, and [v3] to the background in the second image. Where appropriate these tokens could be initialized with CLIP text embeddings semi-representative of the concept (e.g., “person" for both [v1] and [v2]) or simply initialized with random vector embeddings. The latent masks are then initialized by averaging cross-attention maps (see left of Step 1) between the newly defined tokens and the corresponding images across randomly selected $50$ diffusion timesteps. The resulting attention maps are subsequently binarized and refined using DenseCRF, resulting in latent binary masks (see right of Step 1). In Step 2, the tokens are re-optimized for a newly sampled timestep with a combination of Mask Diffusion Loss (3) and Cross-Attention Loss (4) [see text for details]. The new tokens are then used to refine cross-attention and the mask. This alternating optimization continues for a fixed number of steps until masks and tokens converge.
@@ 4 Experimental Results
Table 1: Quantitative evaluation in terms of image similarity and diversity on COCO dataset.
@@
Table 2: Quantitative evaluation based on user study.
@@
  • Similarity of the generated images: We use CLIP [42] and Cosine Similarity [32] scores to evaluate the resemblance of the generated images to the target images in terms of semantic similarity. Given a generated image, we use a pre-trained object detector (e.g. YOLOv8 [24]) to detect the object of interest (corresponding to the input concept) in the generated image and compare their similarity to the target object from MSCOCO images. Table 1 displays the results for the baseline method and our method using CLIP and cosine similarity.
  • Diversity between the generated images: We also consider how diverse the generated images for visual concept-driven generation are for a given prompt by comparing them using LPIPS [58] across different samples in the Table 1.
In Table 1, we present results on the COCO dataset [34]. We randomly selected 120 concepts from 11 COCO object classes, and generated them in various settings using 20 prompts each, creating 8 images for each prompt. We consider our approach with and without data augmentation. Data augmentation makes the generated images more diverse, but it can lead to lower similarity scores when compared to the target images. To account for this, similar to the baseline, we also evaluate our method without data augmentation and calculate semantic similarity using CLIP and cosine similarity scores. Our method without augmentation still performs better in diversity, due to EM-like optimization, allowing it to generate a wider variety of images compared to the baseline. This shows the importance of mask optimization in the process of generating masks.
@@
Figure 6: Qualitative results demonstrating the interaction between different concepts. Here we present the interaction between newly learned concepts and the pre-existing concepts already familiar to the stable diffusion model.
These images were assessed for faithfulness to the target instance and the prompt condition (representing instances, contexts, and interactions) used to generate the images. As shown in Table 2, our method was preferred 71.8% of the time and was considered on par 17.29%. In addition, we conducted another, much larger scale one, using COCO in the same setting as in Tab. 1. In a user study with 17 participants, our approach was chosen 59.44% of the time, compared to the baseline. This clearly illustrates the preference for results obtained using the joint token-mask optimization we propose.
Figure 7: Diverse images for the same prompt. Given input concepts, our method demonstrates the ability to generate a variety of diverse outputs for a single prompt.
@@
Figure 11: Qualitative ablation by removing different components.
Additional Results. In Figure 6, we present the comparative results to show the interaction between newly introduced concepts and the pre-existing ones within the stable diffusion model. Our proposed method is not only able to interact between multiple concepts but also shows its capability to produce diverse outputs for a single prompt, as illustrated in Figure 7.
Furthermore, we demonstrate (in Figure 8) that our approach can interact within a cartoon world context as well. In this scenario, the pre-existing characters (e.g. Obama and Trump) are also depicted in a cartoon-like appearance. Besides, apart from lexical and cartoon concepts, we conduct experiments utilizing real-world concepts and show that our approach can generate multi-concept interactions between pre-existing and new concepts. We compare the qualitative results of our method in Figure 9. The comparative images are taken directly from Break-a-Scene [4]. We can see that our method can generate a better-optimized mask without any spatial supervision. Additionally, the generated images are more aligned with specific subjects and adhere more closely to the provided text prompts.

### v3

@@ Abstract
Index Terms:
personalization, DenseCRF, EM-like Optimization, latent diffusion model, cross-attention.
I Introduction
@@ II Related Work
III Methodology
Figure 2: Overview of our EM-style Optimization Framework. Unique tokens are first assigned to the concepts. In this example, three tokens are assigned: [v1] to the male subject in the first image, [v2] to the female subject in the second image, and [v3] to the background in the second image. Where appropriate these tokens could be initialized with CLIP text embeddings semi-representative of the concept (e.g., “person” for both [v1] and [v2]) or simply initialized with random vector embeddings. The latent masks are then initialized by averaging cross-attention maps (see left of Step 1) between the newly defined tokens and the corresponding images across randomly selected $50$ diffusion timesteps. The resulting attention maps are subsequently binarized and refined using DenseCRF, resulting in latent binary masks (see right of Step 1). In Step 2, the tokens are re-optimized for a newly sampled timestep with a combination of Mask Diffusion Loss (3) and Cross-Attention Loss (4) [see text for details]. The new tokens are then used to refine cross-attention and the mask. This alternating optimization continues for a fixed number of steps until masks and tokens converge.
@@ IV Experimental Results
Figure 9: Qualitative ablation by removing different components.
@@
TABLE I: Quantitative evaluation in terms of image similarity and diversity on COCO dataset.
@@
TABLE II: Quantitative evaluation based on user study.
@@
  • Similarity of the generated images: We use CLIP [41] and Cosine Similarity [32] scores to evaluate the resemblance of the generated images to the target images in terms of semantic similarity. Given a generated image, we use a pre-trained object detector (e.g. YOLOv8 [24]) to detect the object of interest (corresponding to the input concept) in the generated image and compare their similarity to the target object from MSCOCO images. Table I displays the results for the baseline method and our method using CLIP and cosine similarity.
  • Diversity between the generated images: We also consider how diverse the generated images for visual concept-driven generation are for a given prompt by comparing them using LPIPS [58] across different samples in the Table I.
In Table I, we present results on the COCO dataset [34]. We randomly selected 120 concepts from 11 COCO object classes, and generated them in various settings using 20 prompts each, creating 8 images for each prompt. We consider our approach with and without data augmentation. Data augmentation makes the generated images more diverse, but it can lead to lower similarity scores when compared to the target images. To account for this, similar to the baseline, we also evaluate our method without data augmentation and calculate semantic similarity using CLIP and cosine similarity scores. Our method without augmentation still performs better in diversity, due to EM-like optimization, allowing it to generate a wider variety of images compared to the baseline. This shows the importance of mask optimization in the process of generating masks.
@@
These images were assessed for faithfulness to the target instance and the prompt condition (representing instances, contexts, and interactions) used to generate the images. As shown in Table II, our method was preferred 71.8% of the time and was considered on par 17.29%. In addition, we conducted another, much larger scale one, using COCO in the same setting as in Tab. I. In a user study with 17 participants, our approach was chosen 59.44% of the time, compared to the baseline. This clearly illustrates the preference for results obtained using the joint token-mask optimization we propose.
@@
Additional Results. we demonstrate (in Figure 6) that our approach can interact within a cartoon world context as well. In this scenario, the pre-existing characters (e.g. Obama and Trump) are also depicted in a cartoon-like appearance. Besides, apart from lexical and cartoon concepts, we conduct experiments utilizing real-world concepts and show that our approach can generate multi-concept interactions between pre-existing and new concepts. We compare the qualitative results of our method in Figure 7. The comparative images are taken directly from Break-a-Scene [4]. We can see that our method can generate a better-optimized mask without any spatial supervision. Additionally, the generated images are more aligned with specific subjects and adhere more closely to the provided text prompts.